f={(1,2),(2,3),(3,3)},g={(1,0)(2,3),(3,2),(4,8)}
a.Find f+g=

Answer by greenestamps(13195)



(f+g)(1) = f(1)+g(1) = 2+0 = 2
(f+g)(2) = f(2)+g(2) = 3+3 = 6
(f+g)(3) = f(3)+g(3) = 3+2 = 5

Although g(4) is defined, f(4) is not; therefore (f+g)(4) is not defined.

ANSWER: (f+g) = {(1,2),(2,6),(3,5)}